1. An object is 70 µm long and 47.66 µm wide. How long and wide is the object in km? (4)

A) 70 000 km long and 47622 km wide.         

B) 7.0 x 10-8 km long and 4.8 km wide 

C) 7.0 x 10-8 km long and 4.766 x 10-8 km wide

D) 0.7 x 10-8 km long and 4.766 x 10-7 km wide 


2. Which one of the following is the longest length? (4)

A) 100 decimeters C) 105 millimeters E) 10-7 nanometers

B) 10-2 centimeters D) 101 meters



3. A typical E.coli bacteria is about 5 microns (or micrometers) in length. How long is this length in millimeter? (3)

A) 5.000 mm C) 0.005 mm E) 50 mm

B) 5000 mm D) 0.05 mm


4. Which one of the following equations are dimensionally consistent? (3)

A) v = at2 C) v =  

B) v = at2 D) v2 = 2ax

5. How many significant figures does the number 0.000054 have? (2)

A) 4 C) 5

B) 6 D) 2


6. How many significant figures does the number 3.001 x 105 have? (2)

A) 4 C) 7

B) 10 D) 2
































1
Expert's answer
2021-06-07T09:34:13-0400

1. Converting to km, have "70\mu m = 70\times 10^{-6}m = 70\times 10^{-9}km" and "47.66\mu m = 47.66\times 10^{-6}m = 47.66\times 10^{-9}km = 4.766\times 10^{-8}km"

Thus, the answer is C.


2. D


3. 5 microns is "5\times 10^{-6}m = 5\times 10^{-3}mm = 0.005mm". Thus, the answer is C.


4. If a denotes the acceleration, then A and B are consistent.


5. Ignoring all zeros at the beginning, obtain two significant figures (5 and 4). Thus, the answer is D.


6. A.